# Searches for Compact Binary Coalescence Events Using Neural Networks in LIGO/Virgo Third Observation Period

### v1

@@ Abstract
We present the results on the search for the coalescence of compact binary mergers using convolutional neural networks and the LIGO/Virgo data for the O3 observation period. Two-dimensional images in time and frequency are used as input. The analysis is performed in three separate mass regions covering the range for the masses in the binary system from 0.2 to 100 , excluding very asymmetric mass configurations. We explore neural networks trained with input information from pairs of interferometers or all three interferometers together, concluding that the use of the maximum information available leads to an improved performance. A scan over the O3 data set, using the convolutional neural networks, is performed with different fake rate thresholds for claiming detection of at most one event per year or at most one event per week. The latter would correspond to a loose online selection still leading to affordable fake alarm rates. The efficiency of the neutral networks to detect the O3 catalog events is discussed. In the case of a fake rate threshold of at most one event per week, the scan leads to the detection of about of the O3 catalog events. Once the search is limited to the catalog events within the mass range used for neural networks training, the detection efficiency increases up to 70. A further improvement in the search efficiency, using the same kind of algorithms, will require the implementation of new criteria for the suppression of detector glitches.
@@ 1 Introduction
In this paper, we follow closely the analysis procedure in Refs. [5, 6] to search for CBC events in the data from the third LIGO-Virgo-KAGRA observation run (O3). Compared to Ref. [5], which analysed the O2 data, a number of improvements are implemented. New signal regions in the masses of the binary system, covering the whole mass range between 0.2 and 100 , are introduced, and a new approach in determining the CNN working point for signal discrimination, based on the computed false alarm rate (FAR) for the selected candidates, is employed.
@@ 2 Data preparation
The study uses the O3 data [11] from LIGO-Livingston (L1), LIGO-Hanford (H1) and Virgo (V1) interferometers with 4096 Hz sampling rate. The data cover from 1 April 2019 1500 UTC to 27 March 2020 1700 UTC, divided in two periods, denoted as O3a and O3b, separated by a commissioning period of one month in October 2019. This constitutes a total of 155 days of H1-L1-V1 combined data, after applying data quality requirements. A fraction of the data is used to construct background and background plus injected signal images for the purpose of training the CNNs. An additional veto of times containing O3 GW events, as indicated from the GWTC-3 catalog, is included to avoid contamination of the background sample. The signals are generated using the PyCBC package [12, 13, 14] and the waveform approximant IMRPhenomPv2 [15]. Signals are combined with the background data from the different interferometers after taking into account the proper relative orientations, times of arrival and antenna factors.
As presented in Fig. 1, the mass range of the binary system is separated in three regions with increasing masses of the binary components and , with . As in Ref. [5], the low-mass region covers the mass range between 0.2 - 5.0 and a corresponding luminosity distance, , is limited to 100 Mpc; and the high-mass region includes the mass range between 25 and 100 with between 100 Mpc to 1400 Mpc. This is now complemented with an intermediate-mass region covering the mass range between 5 and 25 and in the range 1 to 1000 Mpc. The different limitations in take into account the O3 observed distribution in the data and the expected sensitivity at very large distances. In the CNN training process, the configurations with are excluded. The latter corresponds to very asymmetric mass configurations for which a dedicated CNN search [6] has been performed separately. Figure 1 also collects the masses corresponding to the GWTC-3 catalog events and indicates whether the events are finally detected by this work, as discussed below.
The rest of the CBC event parameters are homogeneously distributed, as indicated in Table 1. This results into signal grids with different generated signals in each mass region.
@@
In order to control the duration of the signals, a low frequency threshold of 80 Hz is applied to the signals in the low- and intermediate-mass regions. For the high-mass region, this threshold is reduced to 25 Hz. The signals duration is limited to five seconds counting backwards from the merger time to remove low frequency components that might confuse the neural network. The generated signals are randomly placed within five second windows of data from each interferometer before being processed. Background and background plus signal segments are whitened following the same prescription as in Ref. [16]. The whitened segments are used to produce spectrograms using -transforms [17] with 400 bins in time and 100 bins in frequency. Finally, the images are processed such that their content has zero average and variance equal to one.
@@ 3 Neural network definition and training
As in previous studies, we adopted a deep CNN ResNet-50 with a 50-layer architecture, as described in Ref. [18]. Binary cross-entropy was used as a loss function along with Adam as the optimizer [19]. The CNN training was based on the first O3a data period. No attempt was made to carry out a new training process for the O3b data since the sensitivity of the instruments remained stable during the whole O3 observation period [20, 21]. A total of four CNNs per mass range were trained covering all the combinations of interferometer inputs: H1-L1, H1-V1, L1-V1 and the triple combination, H1-L1-V1. The use of single interferometers as input to the CNNs was discarded due to the lower performance already observed in the past compared to the use of multiple inputs. Each neural network is trained for up to 12 epochs and the one with the lowest error over the validation set was selected. In all cases the training showed a stable behaviour. A further improvement of the global sensitivity is achieved by combining the outputs of the separate CNNs into a global discriminant. Such combination provides an additional tool for suppressing glitches in the data affecting independently the interferometers and in different time stamps. As in the case of Ref. [6], a simple average of the H1-L1-V1, H1-L1, L1- V1, and H1-V1 CNN outputs has been considered. As shown in Fig. 2, this translates into a better separation of the background and signal and a decrease in the amount of false positives detected by the neural networks.
@@
The performance of the neural networks is presented in Fig. 3 in terms of the receiver operating characteristic (ROC) curves for the separate CNNs and their combination, representing the true positive (TP) versus the false positive (FP) rates. The best performance is achieved by the high mass neural network, reaching high TP values at low FP. This is to a large extent expected since the high-mass signals are generally louder and more visible in the instruments.
@@
As already mentioned, for each image the CNN discriminant is associated to a FAR value, and the claim for observing a CBC candidate is determined by a predefined FAR threshold. The FAR is estimated as a function of the discriminant of the CNN and is defined as , where is the CNN discriminant, is the number of events with a CNN discriminant above or equal to and the period of time analysed. In order to effectively increase the time considered in the FAR calculation, thus reaching very low FAR values, the time slide technique [22] is used. This allows accumulating images (of 5 s of duration each) and accessing FAR values down to years. Using this method, we will consider a possible CBC detection when the combined CNN discrimination has an associated FAR value lower than either one event per year or one event per week, where the latter is used to explore a looser selection more adequate for an online implementation of the algorithm, while maintaining the fake rate at a tenable level.
@@ 5 Results
The CNN global discriminating outputs in the different mass ranges, defined as the average of the corresponding H1-L1-V1, H1-L1, L1-V1, and H1-V1 CNN outputs, are used to search for CBC signals. Here we limit ourselves to the analysis of the data for which all the three interferometers were declared in science mode. A slicing window of five seconds duration was used in steps of 2.5 seconds (leading to a 50 overlap between consecutive images) for each of the interferometers. A scan over the data using different global discriminating values in the range between 0 and 1 is performed. In each case, the corresponding FAR is computed. Figure 6 shows the resulting inverse FAR (IFAR) distributions in the separate mass ranges, in units of years, compared to the expected yields of noise events following Poisson probability distributions.
In the high-mass range an excess of detections above the only-noise prediction is observed. This excess is attributed to the detection of CBC events present in the LIGO/Virgo catalog. It vanishes once the CBC events are excluded from the data. No excess of events above the only-noise prediction is observed in the other mass regions.
Only 51 out of the 90 events from the LIGO/Virgo catalog are inside the part of the data analysed here, with all the three interferometers in science mode. Our CNNs, using a FAR threshold of 1 event per week, detect 26 events, corresponding to an efficiency of about . However, this efficiency also includes CBC events out of the mass range used during the CNN training. Out of the 35 catalog events with chirp masses within our CNN training range, 24 events are detected, corresponding to a detection efficiency of , close to the value anticipated by the signal injection studies.
Figure 7 presents the correlation between the network SNR and the chirp mass. As anticipated, the CNN detections are clustered at large SNR and large masses. Finally, A collects the numerical results for the subset of O3 CBC events for which all the three detectors were in science mode. Improving the detection efficiency for a given FAR rate beyond the observed results using CNNs and two-dimensional images would require the suppression of glitches using dedicated algorithms, beyond the scope of this study.
@@ 6 Summary
We present an update on the search for the coalescence of compact binary mergers using LIGO/Virgo data and convolutional neural networks based on two-dimensional images in time and frequency as input. The analysis is performed in three separate mass regions covering the mass range 0.2 to 100 . A scan over the O3 LIGO/Virgo data set, using a fake rate threshold for claiming detection of at most one event per week, leads to the detection of about of the O3 catalog events. Once the search is limited to the catalog events within the mass range used for neural networks training, the detection efficiency increases up to about 70. A further improvement in the search efficiency, using the same algorithms, will require the implementation of new criteria for the suppression of detector glitches.
@@ Acknowledgements
This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation. The authors also gratefully acknowledge the support of the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO 600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. The authors gratefully acknowledge the Italian Istituto Nazionale di Fisica Nucleare (INFN), the French Centre National de la Recherche Scientifique (CNRS) and the Netherlands Organization for Scientific Research (NWO), for the construction and operation of the Virgo detector and the creation and support of the EGO consortium. The authors thankfully acknowledge the computer resources at MinoTauro and the technical support provided by Barcelona Supercomputing Center (RES-FI-2021-3-0020). This paper has been given LIGO DCC number LIGO-XXXXX. This work is partially supported by the Spanish MCIN/AEI/ 10.13039/501100011033 under the grants SEV-2016-0588, PGC2018-101858-B-I00, and PID2020-113701GB-I00 some of which include ERDF funds from the European Union. IFAE is partially funded by the CERCA program of the Generalitat de Catalunya. M. A-C. is supported by the 2022 FI-00335 grant. This work was carried out within the framework of the EU COST action No. CA17137.
@@ Appendix A CNN results
Table 3 collects the obtained CNN discriminants and the corresponding FAR values for the O3 events. Only the 51 events with the three detectors in science mode have been considered. Results are provided for all the three CNNs corresponding to high-mass, intermediate-mass, and low-mass ranges, as discussed in the body of the text.

### v3

@@ Abstract
We present the results on the search for the coalescence of compact binary mergers using convolutional neural networks and the LIGO/Virgo data for the O3 observation period. Two-dimensional images in time and frequency are used as input. The analysis is performed in three separate mass regions covering the range for the masses in the binary system from 0.2 to 100 , excluding very asymmetric mass configurations. We explore neural networks trained with input information from pairs of interferometers or all three interferometers together, concluding that the use of the maximum information available leads to an improved performance. A scan over the O3 data set, using the convolutional neural networks, is performed with different false rate thresholds for claiming detection of at most one event per year or at most one event per week. The latter would correspond to a loose online selection still leading to affordable false alarm rates. The efficiency of the neutral networks to detect the O3 catalog events is discussed. In the case of a false rate threshold of at most one event per week, the scan leads to the detection of about of the O3 catalog events. Once the search is limited to the catalog events within the mass range used for neural networks training, the detection efficiency increases up to 70. Further improvement in search efficiency, using the same type of algorithms, will require the implementation of new criteria to suppress the remaining major background sources.
@@ 1 Introduction
In this paper, we follow closely the analysis procedure in [5, 6] to search for CBC events in the data from the third LIGO-Virgo-KAGRA observation run (O3). Compared to Ref. [5], which analysed the O2 data, a number of improvements are implemented. New signal regions in the masses of the binary system, covering the whole mass range between 0.2 and 100 , are introduced, and a new approach in determining the CNN working point for signal discrimination, based on the computed false alarm rate (FAR) for the selected candidates, is employed.
@@ 2 Data preparation
The study uses the O3 data [11] from LIGO-Livingston (L1), LIGO-Hanford (H1) and Virgo (V1) interferometers with 4096 Hz sampling rate, obtained from the Gravitational Wave Open Science Center [12]. The data cover from 1 April 2019 1500 UTC to 27 March 2020 1700 UTC, divided in two periods, denoted as O3a and O3b, separated by a commissioning period of one month in October 2019. After applying data quality requirements dealing with the understanding of the interferometer stationary noise budget as well as the identification and suppression of glitches and spectral noise contributions [13, 14], the sample includes a total of 155 days of combined H1-L1-V1 data. A fraction of these data is used to construct background and background plus injected signal images for the purpose of training the CNNs. An additional veto of times containing O3 GW events, as indicated in the GWTC-3 catalog [15, 16, 17], is included to avoid contamination of the background sample. The signals are generated using the PyCBC package [18, 19, 20] and the waveform approximant IMRPhenomPv2 [21]. Signals are combined with the background data from the different interferometers after taking into account the proper relative orientations, times of arrival and antenna factors.
As presented in Figure 1, the mass range of the binary system is separated in three regions with increasing masses of the binary components and , with . As in Ref. [5], the low-mass region covers the mass range between 0.2 - 5.0 and a corresponding luminosity distance, , is limited to 100 Mpc; and the high-mass region includes the mass range between 25 and 100 with between 100 Mpc to 1400 Mpc. This is now complemented with an intermediate-mass region covering the mass range between 5 and 25 and in the range 1 to 1000 Mpc. The different limitations in take into account the O3 observed distribution in the data and the expected sensitivity at very large distances. In the CNN training process, the configurations with are excluded. The latter corresponds to very asymmetric mass configurations for which a dedicated CNN search [6] has been performed separately. Figure 1 also collects the masses corresponding to the GWTC-3 catalog events and indicates whether the events are finally detected by this work, as discussed below. The generated event parameters are uniformly distributed, as indicated in Table 1.
@@
| [M] | [0.17, 4.35] | [4.35, 21.76] | [21.76, 87.06] |
In order to control the duration of the signals, a low frequency threshold of 80 Hz is applied to the signals in the low- and intermediate-mass regions. For the high-mass region, this threshold is reduced to 25 Hz. The signals duration is limited to five seconds counting backwards from the merger time to remove low frequency components that might confuse the neural network. The generated signals are randomly placed within five second windows of data from each interferometer before being processed. Background and background plus signal segments are whitened following the same prescription as in Ref. [22]. The whitened segments are used to produce spectrograms using -transforms [23] with 400 bins in time and 100 bins in frequency. Finally, the images are processed such that their content has zero average and variance equal to one, following the same prescription as in Ref. [5].
@@ 3 Neural network definition and training
As in previous studies, we adopted a deep CNN ResNet-50 with a 50-layer architecture, as described in [24]. Binary cross-entropy was used as a loss function along with Adam as the optimizer [25]. The CNN training was based on the first O3a data period. No attempt was made to carry out a new training process for the O3b data since the sensitivity of the instruments remained stable during the whole O3 observation period. The dataset consists of 250,000 images, of which 115,000 were used for training, 120,000 for testing and 15,000 for validation, evenly distributed into background-only and background with a signal injected. A total of four CNNs per mass range were trained covering all the combinations of interferometer inputs: H1-L1, H1-V1, L1-V1 and the triple combination, H1-L1-V1. The use of a single interferometer as input to the CNNs was discarded due to the lower performance already observed in the past compared to the use of multiple inputs. Each neural network is trained for up to 12 epochs and the one with the lowest error over the validation set was selected. In all cases the training showed a stable behaviour. Keras was used to build and train the CNN with the TensorFlow backend and its implementation on GPUs [26]. Training typically takes two hours per CNN, and overfitting is avoided by checking the CNN performance on the validation set after each epoch. A further improvement of the global sensitivity is achieved by combining the outputs of the separate CNNs into a global discriminant. Such combination provides an additional tool for suppressing glitches in the data affecting independently the interferometers and in different time stamps. As in the case of Ref. [6], a simple average of the H1-L1-V1, H1-L1, L1- V1, and H1-V1 CNN outputs has been considered. As shown in Figure 2, this translates into a better separation of the background and signal and a decrease in the amount of false positives detected by the neural networks.
@@
The performance of the neural networks is presented in Figure 3 in terms of the receiver operating characteristic (ROC) curves for the separate CNNs and their combination, representing the true positive (TP) versus the false positive (FP) rates. The best performance is achieved by the high mass neural network, reaching high TP values at low FP. This is to a large extent expected since the high-mass signals are generally louder and more visible in the instruments.
@@
As already mentioned, for each image the CNN discriminant is associated to a FAR value, and the claim for observing a CBC candidate is determined by a predefined FAR threshold. The FAR is estimated as a function of the discriminant of the CNN and is defined as , where is the CNN discriminant, is the number of events with a CNN discriminant above or equal to and the period of time analysed. In order to effectively increase the time considered in the FAR calculation, thus reaching very low FAR values, the time slide technique [27] is used. This allows accumulating images (of 5 s of duration each) and accessing FAR values down to years-1. Using this method, we will consider a possible CBC detection when the combined CNN discrimination has an associated FAR value lower than either one event per year or one event per week, where the latter is used to explore a looser selection more adequate for an online implementation of the algorithm, while maintaining the false rate at a tenable level.
@@ 5 Results
The CNN global discriminating outputs in the different mass ranges, defined as the average of the corresponding H1-L1-V1, H1-L1, L1-V1, and H1-V1 CNN outputs, are used to search for CBC signals. Here we limit ourselves to the analysis of the data for which all the three interferometers were declared in science mode. A slicing window of five seconds duration was used in steps of 2.5 seconds (leading to a 50 overlap between consecutive images) for each of the interferometers. A scan over the data using different global discriminating values in the range between 0 and 1 is performed. In each case, the corresponding FAR is computed. The computation time for the entire O3 analysis has been of the order of 2000 CPU hours per search (on an Intel Xenon CPU E5-2680 v4 2.4 GHz), which is a significant improvement over the CPU time typically required for known matched-filtering pipelines. Figure 6 shows the resulting inverse FAR (IFAR) distributions in the separate mass ranges, in units of years, compared to the expected yields of noise events following Poisson probability distributions.
In the high-mass range an excess of detections above the only-noise prediction is observed. This excess is attributed to the detection of CBC events present in the GWTC-3 catalog. It vanishes once the CBC events are excluded from the data. No excess of events above the only-noise prediction is observed in the other mass regions.
According to the GWTC-3 catalog, the data analysed in this paper include 50 O3 events in which all three interferometers are in science mode. Using a FAR threshold of one event per week the HM CNN detects 26 of those events, corresponding to an efficiency of . However, only 31 of the 50 events have chirp masses within the CNN training range, and of those 22 are detected, corresponding to a detection efficiency of , close to the value anticipated by the signal injection studies.
Figure 7 presents the correlation between the network SNR and the chirp mass. As expected, given the low detection efficiency of the CNNs for low SNR events observed in Figure 4, detections are clustered at large SNR and large masses. With the current model and for large values of the discriminant, small discriminant variations translate into large variations in FAR, and discriminant values around 0.95 are necessary to reduce the FAR to one event per year. Consequently, several events with values very close to the threshold are not detected. Finally, A collects the numerical results for the subset of O3 CBC events for which all the three detectors were in science mode. Improving the detection efficiency for a given FAR rate beyond the observed results, using CNNs and two-dimensional images, would require a full characterisation of the major remaining noise sources and the implementation of new criteria for their suppression before the CNN algorithms are applied, which is beyond the scope of this study.
@@ 6 Summary
We present an update on the search for the coalescence of compact binary mergers using LIGO/Virgo data and convolutional neural networks based on two-dimensional images in time and frequency as input. The analysis is performed in three separate mass regions covering the mass range 0.2 to 100 . A scan over the O3 LIGO/Virgo data set, using a false rate threshold for claiming detection of at most one event per week, leads to the detection of about of the O3 catalog events. Once the search is restricted to the catalog events within the mass range used for neural networks training, the detection efficiency increases up to about 70. The search is limited by a large rate of false positives originating from background. Further improvement in search efficiency, using the same type of algorithms, will require the implementation of new criteria to suppress the remaining major background sources.
@@ Acknowledgements
This research has made use of data or software obtained from the Gravitational Wave Open Science Center (gwosc.org), a service of the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation, as well as the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. KAGRA is supported by Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan Society for the Promotion of Science (JSPS) in Japan; National Research Foundation (NRF) and Ministry of Science and ICT (MSIT) in Korea; Academia Sinica (AS) and National Science and Technology Council (NSTC) in Taiwan. The authors thankfully acknowledge the computational resources provided by PIC and MinoTauro, and the technical support provided by Barcelona Supercomputing Center (RES-FI-2021-3-0020). This paper has been given LIGO DCC number LIGO-P2400028 and Virgo TDS number VIR-0065A-24. This work is partially supported by the Spanish MCIN/AEI/ 10.13039/501100011033 under the grants SEV-2016-0588, PGC2018-101858-B-I00, and PID2020-113701GB-I00 some of which include ERDF funds from the European Union. IFAE is partially funded by the CERCA program of the Generalitat de Catalunya. M. A-C. is supported by the 2022 FI-00335 grant. This work was carried out within the framework of the EU COST action No. CA17137.
@@ Appendix A CNN results
Table 3 collects the obtained CNN discriminants and the corresponding FAR values for the O3 events. Only the 50 events with the three detectors in science mode have been considered. Results are provided for all the three CNNs corresponding to high-mass, intermediate-mass, and low-mass ranges, as discussed in the body of the text.